OpenStudy (anonymous):

A pitcher exerts 100.0 N of force on a ball with a power output of 4,500 W. What is the velocity of the ball? 0.02 m/s, 45 m/s, 450 m/s, or 4500 m/s ?? :/ thank you!

OpenStudy (michele_laino):

here we have the subsequent relationship between power P, force F and velocity v: \[P = Fv\]

OpenStudy (michele_laino):

so: \[v = \frac{P}{F}\]

OpenStudy (anonymous):

ok! what do we plug in? :/

OpenStudy (michele_laino):

here is next step: \[v = \frac{P}{F} = \frac{{4500}}{{100}}\]

OpenStudy (anonymous):

oh! so we get 45? is 45 m/s our solution?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yay!! thank you:)

OpenStudy (michele_laino):

:)